OpenStudy (anonymous):

Center City East Parking Garage has a capacity of 251 cars more than Center City West Parking Garage. If the combined capacity for the two garages is 1221 cars, find the capacity for each garage. A. Center City East: 485 cars Center City West: 736 cars B. Center City East: 736 cars Center City West: 485 cars C. Center City East: 746 cars Center City West: 475 cars D. Center City East: 475 cars Center City West: 746 cars

OpenStudy (anonymous):

The Capacity of Center City West Parking = x The Capacity of Center City East Parking = x + 251 Parking Capacity of Both Centres = Total Parking Capacity x+x+251=1221 2x+251=1221 2x=1221-251 2x=970 2x/2=970/2 x=485 The Capacity of Center City West Parking = x = 485 The Capacity of Center City East Parking = x + 251 = 485+251=736

OpenStudy (anonymous):

485+736=1221 1221=1221